OpenStudy (anonymous):

30x+8y= -586 -48y+24x=-360 solve for x and y

OpenStudy (anonymous):

Let us first take out a variable i wanna take out y. Why y? Well it is easiest to find the GCF for 8 and 48 (not entirely sure that is the actually acronym GCF I mean but doesnt matter) GCF is 48 so lets make the top equation -48y we do that by multiplying by 6 180x + 48y = -3516 24x - 48y = -360 204 x = -3876 x = -19 now substitute to find y Imma choose the top function 30(-19) +8y = -586 8y = -16 y=-2